OpenStudy (anonymous):

Sophie drove 225 mi south through the Smoky Mountains. Snow during her return trip made her average speed on the return trip 30 mph slower. The total driving time was 8 hr. Find Sophie’s average sped on each part of the trip.

OpenStudy (anonymous):

Let r be the speed to the Smoky Mountains. r t = d, t = d/r Solve the following for r:\[\frac{225}{r}+\frac{225}{r-30}=8 \]r = 45/4 mph or 75 mph .\[\frac{225}{r}+\frac{225}{r}=8 \]With no delay: r = 56.25 mph The initial speed had to be greater than 56.25 mph when there was a delay. Therefore r has to be 75mph on the trip to the Mountains and 45 mph on the return.

OpenStudy (anonymous):

You are a life saver....thank you soooo very much!

OpenStudy (anonymous):

Your welcome.

OpenStudy (madhu.mukherjee.946):

@robtobey nice job

OpenStudy (anonymous):

Thank you and thank you for medal.

OpenStudy (madhu.mukherjee.946):

your welcome